OpenStudy (chihiroasleaf):

calculate the value : \[\sqrt{1+\frac{ 1 }{ 1^2 }+\frac{ 1 }{ 2^2 }} + \sqrt{1+\frac{ 1 }{ 2^2 }+\frac{ 1 }{ 3^2 }}+\sqrt{1+\frac{ 1 }{ 3^2 }+\frac{ 1 }{ 4^2 }}+...+\sqrt{1+\frac{ 1 }{ 2004^2 }+\frac{ 1 }{ 2005^2 }}\]

OpenStudy (anonymous):

That's one sweet sum :P I'm doing.

OpenStudy (anonymous):

Now I have to start over, stupid crash

OpenStudy (freckles):

I was able to find the sum by finding a pattern in the first few sums

OpenStudy (anonymous):

Nope, I got the eqation wrong again, darn

OpenStudy (freckles):

that is from 1 to 2 we have 8/3 and from 1 to 3 we have 15/4 and from 1 to 4 we have 24/5 and so on...

OpenStudy (anonymous):

the trouble with this sum is , it is under roots I'm still thinking what to do

OpenStudy (freckles):

\[\sum_{i=1}^{n}\sqrt{1+\frac{1}{x^2}+\frac{1}{(x+1)^2}}=\frac{n(n+2)}{n+1}\]

OpenStudy (freckles):

that is the pattern I observe

OpenStudy (freckles):

from the first few sums

OpenStudy (rational):

\[\large \sqrt{1+\dfrac{1}{n^2} +\frac{1}{(n+1)^2}} = \sqrt{\dfrac{(n^2+n+1)^2}{n^2(n+1)^2}}\]

OpenStudy (freckles):

also replace x with i *

OpenStudy (freckles):

that was a type-o

OpenStudy (anonymous):

@rational I dont think that equation would solve to the answer.

OpenStudy (rational):

cancel the radical and see if it telescopes

OpenStudy (freckles):

I like @rational 's way more than mine i like telescoping thingys

OpenStudy (anonymous):

@chihiroasleaf Tel me this, is it equated to anything on the left hand side or purely asked for a left hand side value?

OpenStudy (chihiroasleaf):

thanks a bunch @rational :) so, \[\sqrt{1+\frac{ 1 }{ n^2 }+\frac{ 1 }{ (n+1)^2 }} = 1 + \frac{ 1 }{ n(n+1) }\] therefore: \[\sqrt{1+\frac{ 1 }{ 1^2 }+\frac{ 1 }{ 2^2 }} + \sqrt{1+\frac{ 1 }{ 2^2 }+\frac{ 1 }{ 3^2 }}+\sqrt{1+\frac{ 1 }{ 3^2 }+\frac{ 1 }{ 4^2 }}+...+\sqrt{1+\frac{ 1 }{ 2004^2 }+\frac{ 1 }{ 2005^2 }}\] \[=1+\frac{ 1 }{ 1 \times 2 }+1+ \frac{ 1 }{ 2 \times 3 }+.....+1+\frac{ 1 }{2004 \times 2005 }\] \[=2004+\frac{ 1 }{ 1 }-\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }-\frac{ 1 }{ 3 }+....+\frac{ 1 }{ 2004 }-\frac{ 1 }{ 2005 }\] \[=2004+\frac{ 2004 }{ 2005 } =2004\frac{ 2004 }{ 2005} \]

OpenStudy (chihiroasleaf):

@srinuscooby it's purely ask the value of the expression

OpenStudy (rational):

Looks neat !!

OpenStudy (rational):

In general : \[\large \sum\limits_{r=1}^{n} \left[f(r) - f(r+1) \right]= f(1) - f(n+1) \]